A gardener pushes a lawnmower of mass 18kg at a constant speed. To do this requires a force P of 80N directed along the handle, which is at an angle of 40°to the horizontal.


a) Calculate the horizontal frictional force


b) If this frictional force were constant, what force, applied along the handle, would accelerate the mower from rest to 1.2m/s in 2.0s?

1
Expert's answer
2022-01-23T16:17:11-0500

a) "f=P\cos\theta=61.3\text{ N}."

b) By Newton's second law:


"F=\dfrac{ma+f}{\cos\theta}=\dfrac 1{\cos\theta}\bigg(\dfrac{vm}{t}+f\bigg),\\\space\\
F=94\text{ N}."